If x tan 30°+ y cot 60° = 0 and 2x-y tan 45°=1 then find the value of x and y

Grade:11

1 Answers

Shashank Singh
6 years ago
x/v3 + y/v3=0
=> x= -y........ (1)
2x-y=1
2x=1+y............ (2)
putting (1) in (2), we get
-2y= 1+y
y= -1/3
so, x=1/3........ [Ans]